emmahelms:

The perimeter of a rectangle is 16 inches. The equation that represents the perimeter of the rectangle is 2 l plus 2 w equals 16, where l represents the length of the rectangle and w represents the width of the rectangle. Which value is possible for the length of the rectangle?

Gdeinward:

6 is a possible length. L = 6* 2 = 12 W = 2 * 2 = 4 12 + 4 = 16